OpenStudy (anonymous):

Basic Math Questions: (a^b)(a^b)= ? (a^b)(c^d)= ? , (X^ [1/3)])^2 = ?

OpenStudy (anonymous):

a^2b a^b x c^d x^2/3

OpenStudy (callisto):

\[ a^{b} a ^{b} = a ^{2b}\]\[a ^{b}c ^{d}= a ^{b}c ^{d}\]\[(x ^{1/3})^{2} = x ^{2/3}\]

OpenStudy (anonymous):

for the third one, are we adding the 1/3 s? and in the first one, is that like adding the powers together?

OpenStudy (callisto):

first one = adding power third multiplying the power together

OpenStudy (anonymous):

nope, 1/3 + 2 = 4/3

OpenStudy (callisto):

1/3 + 2 = 7/3 .... and it's not adding the power!!!

OpenStudy (anonymous):

is this how we get x^ (2/3)? (1/3)+(1/3) = (2/3) ?

OpenStudy (callisto):

no! (1/3)*2 = (2/3)

OpenStudy (callisto):

\[ x ^{a}\times x ^{b} = x ^{a+b}\]\[(x ^{a})^{b} = x ^{ab}\]

OpenStudy (anonymous):

Perfect! Thanks :)

OpenStudy (callisto):

Welcome :)